# A Frank-Wolfe Algorithm for Oracle-based Robust Optimization

### v1

@@ 1 Introduction
This approach was applied e.g. in Bettiol et al. [2023], Al Dahik et al. [2020]. Problem (2) is of the form (RO); however, since $\mathcal{Z}$ can represent any combinatorial problem, a polynomial-sized description of the set $\operatorname*{conv}\left(\mathcal{Z}\right)$ may be unknown or does not exist, rendering the dualization approach intractable. On the other hand, linearly optimizing over $\operatorname*{conv}\left(\mathcal{Z}\right)$ is equivalent to linearly optimizing over $\mathcal{Z}$, which can often be done efficiently by problem-specific algorithms.
@@
All the above examples motivate the use of so-called oracle-based algorithms, where we assume that the feasible set $\mathcal{X}$ can only be accessed by a linear optimization oracle. More precisely, a linear optimization oracle is an algorithm which returns for every $c\in\mathcal{U}$ an optimal solution of the deterministic problem $\min_{x\in\mathcal{X}}\ c^{\top}x$, alleviating the need for a polynomial-sized formulation of $\mathcal{X}$. Instead, any linear optimization algorithm over $\mathcal{X}$ can be used. Note, that linear optimization over sets of the form $\mathcal{X}=\operatorname*{conv}\left(\mathcal{Z}\right)$ (appearing in the examples above) can be performed by linear optimization over the combinatorial set $\mathcal{Z}$ itself. Hence combinatorial algorithms for the underlying deterministic problem can be used.
@@ 1.1.3 Oracle-based Robust Optimization
Finally, in Bettiol et al. [2023] a simplicial decomposition algorithm is presented, which can be compared to a sub-gradient version of the the Fully-Corrective Frank-Wolfe algorithm. Iteratively, the algorithm uses a linear minimization oracle over $\mathcal{X}$ and a convex hull oracle which returns an optimal solution to the problem
@@
Oracle-based algorithms were also used to solve two-stage robust binary optimization problems [Kämmerling and Kurtz, 2020], min-max-min robust combinatorial optimization problems [Buchheim and Kurtz, 2017], and for robust planning of production routing [Borumand et al., 2024]. In Buchheim [2020], the author proves that in general no oracle-polynomial algorithms exist (unless P=NP) for Problem (RO) if $\mathcal{X}$ is binary.
@@ 2.1 Objective Smoothing
       $\mathbf{v}_{t}=\operatorname*{arg\,min}\limits_{\mathbf{v}\in\mathcal{X}}\langle\mathbf{v},\nabla f_{\mu}(\mathbf{x}_{t})\rangle$
@@ 2.2 Oracle Complexity via Frank-Wolfe Algorithms
       $\mathbf{v}_{t}=\operatorname*{arg\,min}\limits_{\mathbf{v}\in\mathcal{X}}\langle\mathbf{v},\nabla f_{\mu_{t}}(\mathbf{x}_{t})\rangle$
@@ 2.3 Function and Gradient Evaluations
The inner subproblem thus amounts to an Euclidean projection onto the uncertainty set. We focus on budgeted uncertainty of the form
@@
Projecting on the convex hull of a finite scenario set $\{c_{s}\}_{s\in S}$ amounts to solving the quadratic problem:
@@
Budgeted uncertainty is captured by a continuous knapsack polytope. Dualizing the budget constraint makes the projection problem fully separable; this property is leveraged in particular by breakpoint algorithms for diagonal quadratic convex continuous knapsack problems. We refer the reader to Patriksson and Strömberg [2015] for a recent review of solution approaches, from which we implement the two-pegging breakpoint algorithm.

### v2

@@ 1 Introduction
This approach was applied e.g. in Bettiol et al. [2023], Al Dahik et al. [2020]. Problem (2) is of the form (RO); however, since $\mathcal{Z}$ can represent any combinatorial problem, a polynomial-sized description of the set $\operatorname*{conv}\left(\mathcal{Z}\right)$ may be unknown or does not exist, rendering the dualization approach intractable. On the other hand, linearly optimizing over $\operatorname*{conv}\left(\mathcal{Z}\right)$ is equivalent to linearly optimizing over $\mathcal{Z}$, which can often be done by problem-specific algorithms.
@@
All the above examples motivate the use of so-called oracle-based algorithms, where we assume that the feasible set $\mathcal{X}$ can only be accessed by a linear optimization oracle. More precisely, a linear optimization oracle is an algorithm which returns for every $c\in\mathcal{U}$ an optimal solution of the deterministic problem $\min_{x\in\mathcal{X}}\ c^{\top}x$, alleviating the need for a polynomial-sized formulation of $\mathcal{X}$. Instead, any linear optimization algorithm over $\mathcal{X}$ can be used. Hence, combinatorial algorithms for the underlying deterministic problem can be used.
@@ 1.1.3 Oracle-based Robust Optimization
Finally, in Bettiol et al. [2023] a simplicial decomposition algorithm is presented, which can be compared to a sub-gradient version of the Fully-Corrective Frank-Wolfe algorithm. Iteratively, the algorithm uses a linear minimization oracle over $\mathcal{X}$ and a convex hull oracle which returns an optimal solution to the problem
@@
Oracle-based algorithms were also used to solve two-stage robust binary optimization problems [Kämmerling and Kurtz, 2020], min-max-min robust combinatorial optimization problems [Buchheim and Kurtz, 2017], and for robust planning of production routing [Borumand et al., 2024]. In Buchheim [2020], the author proves that in general no oracle-polynomial algorithm exists for Problem (RO) if the uncertainty set contains a finite number of scenarios and $\mathcal{X}$ is binary.
@@ 2.1 Objective Smoothing
       $g_{t}\leftarrow\nabla f_{\mu}(\mathbf{x}_{t})$ (calling the projection oracle on $\mathcal{U}$)
       $\mathbf{v}_{t}=\operatorname*{arg\,min}\limits_{\mathbf{v}\in\mathcal{X}}\langle\mathbf{v},g_{t}\rangle$ (calling the linear optimization oracle over $\mathcal{X}$)
@@ 2.2 Oracle Complexity via Frank-Wolfe Algorithms
       $g_{t}\leftarrow\nabla f_{\mu}(\mathbf{x}_{t})$ (calling the projection oracle on $\mathcal{U}$)
       $\mathbf{v}_{t}=\operatorname*{arg\,min}\limits_{\mathbf{v}\in\mathcal{X}}\langle\mathbf{v},g_{t}\rangle$ (calling the linear optimization oracle over $\mathcal{X}$)
@@ 2.3 Function and Gradient Evaluations
The inner subproblem thus amounts to an Euclidean projection onto the uncertainty set. Hence, for classical budgeted uncertainty of the form
@@
the projection problem results in a quadratic knapsack problem. Dualizing the knapsack constraint makes the projection problem fully separable; this property is leveraged in particular by breakpoint algorithms for diagonal quadratic convex continuous knapsack problems. We refer the reader to Patriksson and Strömberg [2015] for a recent review of solution approaches, from which we implement the two-pegging breakpoint algorithm.
In case the uncertainty set is the convex hull of a finite scenario set $\{c_{s}\}_{s\in S}$, projecting on $\mathcal{U}$ amounts to solving the quadratic problem: